In a collision of sufficient force, automobile air bags respond by electrically triggering the explosive decomposition of sodium azide (NaN3) to its elements. A 29.4−g sample of sodium azide was decomposed, and the nitrogen gas generated was collected over water at 26°C. The total pressure was 733.8 mmHg. How many liters of dry N2 were generated?

1
Expert's answer
2021-03-02T01:03:44-0500

2NaN3 = 2Na + 3N2

M(NaN3) - 65 g/mol; n(NaN3) = 29.4/65= 0.452 mol; n(N2) = 3/2 n(NaN3)= 0.678 mol

PV=nRT

1 mm Hg = 133.3 Pa, 733.8 mmHg= 97815.54 Pa= 97.815 kPa;

V = nRT/P = 0.678 * 8.314 * (273.15 + 26) / 97.815 = 17.24 liters of N2